# TempoNet: Slack-Quantized Transformer-Guided Reinforcement Scheduler for Adaptive Deadline-Centric Real-Time Dispatchs

### v1

@@ 4.1 Experimental Setup
We conduct comprehensive evaluations across three computational scenarios: uniprocessor periodic scheduling with standardized task configurations, industrial multi-core workloads with mixed-criticality workflows, and large-scale multiprocessor systems with 100–600 tasks. Our framework is benchmarked against established scheduling approaches including classical schedulers (RM, EDF), feedforward Deep Q-Network (FF-DQN), Dynamic Importance-aware Online Scheduling (DIOS), and quantum-inspired optimization methods. Evaluation Metrics: Performance assessment employs deadline compliance rate, average response time, and computational overhead. Together, these two components form the core of TempoNet’s learning dynamics, as detailed in Sections N and O. As discussed across Sections P, Q, R, and S, these results collectively characterize both the strengths and the operational limits of TempoNet under diverse real-time conditions. As detailed across Sections I, J, K, and L, these studies collectively characterize the impact of slack representation, sparse-attention efficiency, robustness properties, and system-level behavior.
@@ A.1 Definitions and Policy Formalization
Let $\mathbf{s}=[s_{1},\dots,s_{N}]^{\top}$ denote a vector representing the temporal laxity (slack) of $N$ active tasks, where each coordinate $s_{i}$ is bounded within the compact interval $0,S_{\max}]$. We consider two distinct families of scheduling policies:
@@ Appendix E Differentiability and complexity of the masked-greedy mapping
where $m$ is the number of cores to fill and $N$ is the number of available tasks.
@@ Appendix F Scheduling rationale: slack versus SRPT, and run-level interpretability metrics
Consider two tasks that arrive at time $t=0$ with the following parameters: task $A$ has remaining processing $c_{A}=1$ and deadline $d_{A}=100$; task $B$ has remaining processing $c_{B}=2$ and deadline $d_{B}=2.2$. SRPT schedules the shorter job $A$ first, finishing it by $t=1$, then executes $B$ and completes $B$ at $t=3$, which misses $B$’s deadline. A slack-minimizing policy computes initial slacks $s_{A}=99$ and $s_{B}=0.2$ and therefore schedules $B$ first, completing both tasks before their deadlines. This simple instance generalizes: whenever a job with a slightly larger remaining time has a much earlier deadline, SRPT may prioritize the less urgent job and cause the urgent one to miss its deadline, whereas slack-aware policies avoid this failure mode.
@@ I.2 Sparse-Attention Micro-Benchmark
KS-test $p$-value 0.18 (no bias)
@@
B: Sinkhorn 90.6 12.1 860 +0.5 pp, $\times$2 latency
@@
Figure 9: Real-time boundary: synthetic inference latency versus number of active tasks $N$ with a 1 ms budget line. Vertical dashed line marks the interpolated upper bound $N_{\max}$. Replace with measured timings for an exact per-system bound.
@@
The attention scores are sparsified by retaining only the top-k values per query within predefined blocks. Let the attention score matrix be $A\in\mathbb{R}^{(N+1)\times(N+1)}$, where $N$ is the number of tasks. The matrix is partitioned into blocks of size $B\times B$, where $B$ is the block size determined based on the task count and hardware constraints. For each query vector in a block, we compute the top-k attention scores within its corresponding key block. The sparsified attention matrix $\tilde{A}$ is then given by:
@@
The theoretical complexity of the sparse attention mechanism is $O(N^{1.1})$ on average, achieved through the combination of block Top-k and chunking. To validate this, we measured the wall-clock time for attention computation across task sets of size $N$ ranging from 10 to 600 tasks. The results, plotted in Figure 1, show that the time $T(N)$ fits the model $T(N)=c\cdot N^{1.1}+d$, where $c$ and $d$ are constants determined via linear regression on log-transformed data. The coefficient of determination ($R^{2}$) exceeded 0.98, confirming the scalability. The measurement setup used an NVIDIA V100 GPU with fused batched sparse kernels from the CUDA toolkit, ensuring optimal hardware utilization.
@@
where $\tilde{s}_{i}(t)$ refers to the quantized slack index of task $i$ at time $t$ as per Equation (5) in the main text, representing the task’s urgency level, and $a_{c}(t)\in\{1,2,\ldots,m\}^{N}$ is a core assignment vector where each component $a_{c}^{(i)}(t)$ indicates the core index to which task $i$ is currently assigned, with $m$ being the number of cores and $N$ the total number of tasks.
@@
Consider one self-attention layer applied to a sequence of $N$ tokens. The sequence is partitioned into $m$ non-overlapping blocks of equal size $B$, so that $N=mB$. Inside each block we compute full (dense) attention among the block tokens. For interactions across blocks, each block is summarized (for example by pooled projections or a small set of representatives) and chooses a small set of other blocks to attend to. Denote by $k$ the average number of other blocks selected per block. We count only raw query–key dot products per attention head per layer and ignore projection and constant overheads.
@@
The three terms on the right-hand side correspond respectively to intra-block pairwise scores, the aggregated cost of block selection across all blocks, and inter-block score evaluations incurred when each query inspects up to $kB$ external keys.
@@
If both $B$ and $k$ are constants independent of $N$, the complexity in (86) is linear in $N$. If $B$ and $k$ scale like $\log N$, the cost grows polylogarithmically times $N$. If $B=\Theta(\sqrt{N})$ the dominant contributions scale as $N^{3/2}$ and the per-layer cost becomes super-linear.
@@
Substituting $C=\Theta(N/\log N)$, $B=\Theta(\sqrt{N})$ and $k=\Theta(\sqrt{N})$ and simplifying shows that the intra-chunk term dominates for sufficiently large $N$, yielding the asymptotic estimate
@@
The analytic count in (92) does not account for implementation optimizations. Two effects typically reduce the observed runtime exponent. First, symmetric sparsity patterns inside blocks allow optimized GPU kernels to fuse row- and column-wise accesses, reducing effective memory traffic and lowering constant factors. Second, fitting a power law to wall-clock times over a finite $N$ range, combined with caching and kernel fusion, often produces an apparent exponent smaller than the asymptotic one. Empirically, these practicalities can transform the theoretical $N^{1.5}$ scaling into a measured behaviour close to $N^{1+\varepsilon}$ with small $\varepsilon$; the experiments reported in the paper fitted an exponent near $1.1$ with high goodness-of-fit.
@@
where the meaning of each symbol is as stated above. Fixing $B$ and $k$ keeps $C(N)$ linear in $N$; allowing either to grow with $N$ can push the cost into super-linear regimes.
@@
Measured (optimized CUDA kernel, finite $N$ range) $\Theta(N^{1.1})$ (empirical fit). Observed when kernel fusion and symmetric sparsity reduce constants.
@@
where $I\{\cdot\}$ denotes the indicator function that equals 1 if the enclosed condition is true and 0 otherwise, $c_{i}(t)$ represents the remaining execution time of task $i$ at time $t$, and $d_{i}^{(k)}$ is the absolute deadline of the $k$-th job instance of task $i$. The first term rewards the completion of a job within the current time step, while the second term penalizes a deadline miss for any active job. This design directly encodes the objective of hard real-time systems: to maximize the number of met deadlines and minimize misses, as each missed deadline can lead to system failure or severe degradation in safety-critical applications.
@@
where $N$ is the current number of tokens, $s_{i}$ is the slack of token $i$, and $\mathbf{1}\{\cdot\}$ is the indicator function. In this test we observed that when $p>0.30$ the Top-k kernel load increases markedly and attention statistics indicate over-concentration.

### v2

@@ 4.1 Experimental Setup
We conduct comprehensive evaluations across three computational scenarios: uniprocessor periodic scheduling with standardized task configurations, industrial multi-core workloads with mixed-criticality workflows, and large-scale multiprocessor systems with 100–600 tasks. Our framework is benchmarked against established scheduling approaches including classical schedulers (RM, EDF), feedforward Deep Q-Network ($F$-DQN), Dynamic Importance-aware Online Scheduling (DIOS), and quantum-inspired optimization methods. Evaluation Metrics: Performance assessment employs deadline compliance rate, average response time, and computational overhead. Together, these two components form the core of TempoNet’s learning dynamics, as detailed in Sections N and O. As discussed across Sections P, Q, R, and S, these results collectively characterize both the strengths and the operational limits of TempoNet under diverse real-time conditions. As detailed across Sections I, J, K, and L, these studies collectively characterize the impact of slack representation, sparse-attention efficiency, robustness properties, and system-level behavior.
@@ A.1 Definitions and Policy Formalization
Let $\mathbf{s}=[s_{1},\dots,s_{N}]^{\top}$ denote a vector representing the temporal laxity (slack) of NN active tasks, where each coordinate $s_{i}$ is bounded within the compact interval $0,S_{\max}]$. We consider two distinct families of scheduling policies:
@@ Appendix E Differentiability and complexity of the masked-greedy mapping
where $m$ is the number of cores to fill and NN is the number of available tasks.
@@ Appendix F Scheduling rationale: slack versus SRPT, and run-level interpretability metrics
Consider two tasks that arrive at time $t=0$ with the following parameters: task AA has remaining processing $c_{A}=1$ and deadline $d_{A}=100$; task $B$ has remaining processing $c_{B}=2$ and deadline $d_{B}=2.2$. SRPT schedules the shorter job AA first, finishing it by $t=1$, then executes $B$ and completes $B$ at $t=3$, which misses $B$’s deadline. A slack-minimizing policy computes initial slacks $s_{A}=99$ and $s_{B}=0.2$ and therefore schedules $B$ first, completing both tasks before their deadlines. This simple instance generalizes: whenever a job with a slightly larger remaining time has a much earlier deadline, SRPT may prioritize the less urgent job and cause the urgent one to miss its deadline, whereas slack-aware policies avoid this failure mode.
@@ I.2 Sparse-Attention Micro-Benchmark
KS-test pp-value 0.18 (no bias)
@@
B: Sinkhorn 90.6 12.1 860 +0.5 $p$, $\times$2 latency
@@
Figure 9: Real-time boundary: synthetic inference latency versus number of active tasks NN with a 1 ms budget line. Vertical dashed line marks the interpolated upper bound $N_{\max}$. Replace with measured timings for an exact per-system bound.
@@
The attention scores are sparsified by retaining only the top-k values per query within predefined blocks. Let the attention score matrix be $A\in\mathbb{R}^{(N+1)\times(N+1)}$, where NN is the number of tasks. The matrix is partitioned into blocks of size $B\times B$, where $B$ is the block size determined based on the task count and hardware constraints. For each query vector in a block, we compute the top-k attention scores within its corresponding key block. The sparsified attention matrix $\tilde{A}$ is then given by:
@@
The theoretical complexity of the sparse attention mechanism is $O(N^{1.1})$ on average, achieved through the combination of block Top-k and chunking. To validate this, we measured the wall-clock time for attention computation across task sets of size NN ranging from 10 to 600 tasks. The results, plotted in Figure 1, show that the time $T(N)$ fits the model $T(N)=c\cdot N^{1.1}+d$, where $c$ and $d$ are constants determined via linear regression on log-transformed data. The coefficient of determination ($R^{2}$) exceeded 0.98, confirming the scalability. The measurement setup used an NVIDIA V100 GPU with fused batched sparse kernels from the CUDA toolkit, ensuring optimal hardware utilization.
@@
where $\tilde{s}_{i}(t)$ refers to the quantized slack index of task $i$ at time $t$ as per Equation (5) in the main text, representing the task’s urgency level, and $a_{c}(t)\in\{1,2,\ldots,m\}^{N}$ is a core assignment vector where each component $a_{c}^{(i)}(t)$ indicates the core index to which task $i$ is currently assigned, with $m$ being the number of cores and NN the total number of tasks.
@@
Consider one self-attention layer applied to a sequence of NN tokens. The sequence is partitioned into $m$ non-overlapping blocks of equal size $B$, so that $N=mB$. Inside each block we compute full (dense) attention among the block tokens. For interactions across blocks, each block is summarized (for example by pooled projections or a small set of representatives) and chooses a small set of other blocks to attend to. Denote by $k$ the average number of other blocks selected per block. We count only raw query–key dot products per attention head per layer and ignore projection and constant overheads.
@@
The three terms on the right-hand side correspond respectively to intra-block pairwise scores, the aggregated cost of block selection across all blocks, and inter-block score evaluations incurred when each query inspects up to kBkB external keys.
@@
If both $B$ and $k$ are constants independent of NN, the complexity in (86) is linear in NN. If $B$ and $k$ scale like $\log N$, the cost grows polylogarithmically times NN. If $B=\Theta(\sqrt{N})$ the dominant contributions scale as $N^{3/2}$ and the per-layer cost becomes super-linear.
@@
Substituting $C=\Theta(N/\log N)$, $B=\Theta(\sqrt{N})$ and $k=\Theta(\sqrt{N})$ and simplifying shows that the intra-chunk term dominates for sufficiently large NN, yielding the asymptotic estimate
@@
The analytic count in (92) does not account for implementation optimizations. Two effects typically reduce the observed runtime exponent. First, symmetric sparsity patterns inside blocks allow optimized GPU kernels to fuse row- and column-wise accesses, reducing effective memory traffic and lowering constant factors. Second, fitting a power law to wall-clock times over a finite NN range, combined with caching and kernel fusion, often produces an apparent exponent smaller than the asymptotic one. Empirically, these practicalities can transform the theoretical $N^{1.5}$ scaling into a measured behaviour close to $N^{1+\varepsilon}$ with small $\varepsilon$; the experiments reported in the paper fitted an exponent near $1.1$ with high goodness-of-fit.
@@
where the meaning of each symbol is as stated above. Fixing $B$ and $k$ keeps $C(N)$ linear in NN; allowing either to grow with NN can push the cost into super-linear regimes.
@@
Measured (optimized CUDA kernel, finite NN range) $\Theta(N^{1.1})$ (empirical fit). Observed when kernel fusion and symmetric sparsity reduce constants.
@@
where $I\{\cdot\}$ denotes the indicator function that equals 1 if the enclosed condition is true and 0 otherwise, $c_{i}(t)$ represents the remaining execution time of task $i$ at time $t$, and $d_{i}^{(k)}$ is the absolute deadline of the $k$-th job instance of task ii. The first term rewards the completion of a job within the current time step, while the second term penalizes a deadline miss for any active job. This design directly encodes the objective of hard real-time systems: to maximize the number of met deadlines and minimize misses, as each missed deadline can lead to system failure or severe degradation in safety-critical applications.
@@
where NN is the current number of tokens, $s_{i}$ is the slack of token $i$, and $\mathbf{1}\{\cdot\}$ is the indicator function. In this test we observed that when $p>0.30$ the Top-k kernel load increases markedly and attention statistics indicate over-concentration.